OpenStudy (anonymous):

I have done 4 Bumps already! Help!!!!!!!!!!!!!!!!!!!Medal will be received it work is shown. And if I really like it! You're a fan! Which mathematical sentence represents the solution for d in the equation 6e = ef + 3d? For choices just access the link, if you cannot then just type the answer you got and ill see. Thanks

OpenStudy (anonymous):

Thanks @justin12lo for the medal!

OpenStudy (anonymous):

Can anyone figure this one out?

OpenStudy (anonymous):

cant see the image, says The system is unable to access the cookie that was assigned to your computer upon login. The most likely cause of this problem is an improper access to the course via a bookmark, browser window that was left open, or an improper exit from the system....

OpenStudy (anonymous):

Can you do it without the choices?

OpenStudy (swissgirl):

What do you think? I think its A

OpenStudy (anonymous):

mhmm let me check

OpenStudy (swissgirl):

Ok its def not A :) I was kidding

OpenStudy (e.cociuba):

6e = ef + 3d subtract ef from each side

OpenStudy (swissgirl):

Basically you are solving for d meaning that you gotta isolate d

OpenStudy (e.cociuba):

6e = ef + 3d - ef - ef 6e - ef = ef -ef +3d 6e - ef = 3d divide by 3 2e - ef/3 =d

OpenStudy (e.cociuba):

Do u understand hun?

OpenStudy (swissgirl):

@e.cociuba is correct

OpenStudy (whpalmer4):

Sorry, misread a step!

OpenStudy (e.cociuba):

It's kool :)) No worries! :)

OpenStudy (e.cociuba):

Thanks for the support @swissgirl Woot Woot! U can be my cheerleader! LOL ;D

OpenStudy (swissgirl):

Awwwww yeaaaaaa :)

OpenStudy (e.cociuba):

Lol ;P

OpenStudy (anonymous):

Sorry guys I got banned for making a troll account? But I never did? ;(

OpenStudy (anonymous):

But I'm back not the ban is done :D